# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 7.1 Tree-like grand-coalition-first action models
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
@@
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history $\theta$ from $r$ to $s$.
@@
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$;
@@
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
@@
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
@@
Assume (1). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $a\in\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is an $\mathtt{AG}$-history from $r$ to $r$. Assume $(r,\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{m},r)$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$. Then, $(r,\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{m}|_{a},r)$ is an $a$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$. Let $(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$ be the $a$-history from $r$ to $s$. For every $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{i}_{\mathtt{AG}})$. Then $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
    Let $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ be an $\mathtt{AG}$-history from $r$ to $s$. We claim $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},$ $s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are identical.
    Assume $n\neq m$. Let $a\in\mathtt{AG}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then they are different. We have a contradiction.
    Assume $\sigma^{i}_{\mathtt{AG}}\neq\lambda^{i}_{\mathtt{AG}}$ for some $1\leq i\leq n$. Then there is $a\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{a}\neq\lambda^{i}_{\mathtt{AG}}|_{a}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then, they are different. We have a contradiction.
    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$. Let $a\in\mathtt{AG}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then, they are different. We have a contradiction.
@@
Assume (3). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$. For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{m},s_{m})$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$. Let $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ be the $\mathtt{AG}$-history from $r$ to $s$. Then, $,(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},$$\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is an $\mathrm{C}$-history from $r$ to $s$.
    Assume $(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$ which is different from $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},$ $s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$.
    For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
    Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. We have a contradiction.
@@
It is easy to see that (1) is a special case of (2).
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model. We are done if we can show for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Let $s\in\mathtt{ST}$, $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, there are at least two paths $(s,\sigma_{\mathtt{AG}},w)$ and $(s,\sigma^{\prime}_{\mathtt{AG}},w)$ from $s$ to $w$. Then, there are at least two paths $(r,\dots,s,\sigma_{\mathtt{AG}},w)$ and $(r,\dots,s,\sigma^{\prime}_{\mathtt{AG}},w)$ from r to w. Note that $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like grand-coalition-first action model. Then, there is a contradiction.
It is easy to find a clear grand-coalition-first action model with a loop. Then, it is not tree-like.
@@
In the literature, Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models. They are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined here. It is required in their definition that for every state, there is a unique computation, which is a sequence of states, from the root to the state. Consequently, they are not clear.

### v2

@@ 6.1 Tree-like grand-coalition-first action models
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
@@
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
Note every tree-like model has a unique root.
The following result gives two alternative definitions of tree-like grand-coalition-first action models.
@@
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
@@
Assume (1). Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$. We want to show that there is an unique $\mathtt{AG}$-history from $r$ to $s$. Let $a\in\mathtt{AG}$.
Assume $s=r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$.
Assume there is an $\mathtt{AG}$-history $\theta_{\mathtt{AG}}$ from $r$ to $r$ which is different from $r$. It is impossible that $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. It is easy to check that $\theta_{a}=(s_{0},\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ is an $a$-history from $r$ to $r$. Clearly, $\theta_{a}$ is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $s\neq r$.
Let $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{a}=(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$.
For every $i$ such that $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\sigma^{i}_{\mathtt{AG}})$. Then, $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
Assume there is a different $\mathtt{AG}$-history $\theta^{\prime}_{\mathtt{AG}}$ from $r$ to $s$. Note it is impossible $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathtt{AG}}=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
We want to show $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for all $i$.
Assume $\sigma^{i}_{\mathtt{AG}}\neq\lambda^{i}_{\mathtt{AG}}$ for some $1\leq i\leq n$. Then, there is $b\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{b}\neq\lambda^{i}_{\mathtt{AG}}|_{b}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{b},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{b},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{b},t_{1},\dots,\lambda_{\mathtt{AG}}^{m}|_{b},t_{m})$ are $b$-histories from $r$ to $s$, which are different. We have a contradiction.
@@
Assume (3). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is an unique $\mathrm{C}$-history from $r$ to $s$.
Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$.
Assume there is a $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$ which is different from $r$. Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$.
For every $1\leq i\leq n$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then, $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$. We have a contradiction.
Assume $s\neq r$.
Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. Note it is impossible $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. It is easy to see $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Assume there is a different $\mathrm{C}$-history $\theta^{\prime}_{\mathrm{C}}$ from $r$ to $s$. Again, it is impossible $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathrm{C}}=(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$.
For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then, $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
It is easy to check that $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. We have a contradiction.
@@
(1) is a special case of (2).
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root. Let $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. It suffices to show $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. It is easy to check that both $(\theta_{\mathtt{AG}},\sigma_{\mathtt{AG}},t)$ and $(\theta_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}},t)$ are $\mathtt{AG}$-histories from $r$ to $t$, which are different. We have a contradiction.
It is easy to find a clear grand-coalition-first action model with a loop, which is not tree-like.
@@
Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section. For the former, it is only required that for every state, there is a unique computation, which is a sequence of states, from the root to the state.